OpenStudy (anonymous):

Can someone help me please! I'm on the last question and its killing me! The navigator has to program the ship’s computer so that it will be prepared to intercept the satellite. To program it, the satelite’s path is set equal to the ship’s. The programmed equation is –4x – 2 = –x + 3. Explain to the navigator how the programmed equation can be graphed to find the point of intersection. Use complete sentences.

OpenStudy (anonymous):

@sweetsunray please tell me you understand this and are not as lost as me

OpenStudy (anonymous):

@sourwing do you get this?

OpenStudy (anonymous):

basically you're asked to write down every step you need to take to solve the equation with x= value

OpenStudy (aum):

The programmed equation is –4x – 2 = –x + 3. Plot the straight line y = -4x - 2 Plot the second line y = -x + 3 The point where the two lines intersect is the solution.

OpenStudy (anonymous):

Thanks